OpenStudy (anonymous):

Write the equation of a line that passes through the point (1,-3) and is parallel to the line 7x + y = -5.

OpenStudy (anonymous):

A. y = -7x + 4 B. y = -4x + 7 C. y = 4x + -7 D. y = 28x

OpenStudy (anonymous):

\(y=-7x+b\) \(-7+b=-3\) ==> \(b=4\) so : \(y=-7x+4\)